# Foundation Models for Autonomous Driving System: An Initial Roadmap

### v1

@@ 3. Foundation Models in Vehicle
3.1. FM-Enhanced Autonomous Driving
Existing integration of FMs into autonomous driving systems can be roughly categorized into perception and scene understanding, navigation and planning, decision-making and control, and end-to-end autonomous driving. In this section, we briefly discuss representative techniques and identify the challenges and opportunities. For more background and technique details, we refer readers to prior works (Wu et al., 2024; Gao et al., 2024; Cui et al., 2024c; Zhou et al., 2024b).
3.1.1. Perception and Scene Understanding
FMs enhance perception in autonomous driving by enabling context-aware environmental understanding (Zhou et al., 2024b). VLMs such as LLaVA (Liu et al., 2023b) and GPT-4V (OpenAI, 2023) support tasks like 3D open-vocabulary object detection (Najibi et al., 2023; Pan et al., 2024b), language-guided retrieval (Romero et al., 2023; Wei et al., 2024a), and visual question answering (VQA) (Qian et al., 2024; Choudhary et al., 2024; Nie et al., 2025; Liu et al., 2024). Examples include OpenScene (Peng et al., 2023) for zero-shot 3D semantic segmentation and NuScenes-QA (Qian et al., 2024) for VQA benchmarks. Additionally, DriveVLM (Tian et al., 2024a) employs chain-of-thought reasoning for scene analysis, while DriveDreamer (Wang et al., 2025) predicts future states for proactive responses.
Challenge & Opportunity: Hallucination Mitigation. While FMs (e.g., LLMs and VLMs) have achieved significant advancements in autonomous driving tasks, hallucination remains a critical challenge for real-world deployment. FMs are particularly prone to hallucination, which refers to generating outputs that are factually incorrect, inconsistent, or nonsensical (Sun et al., 2025; Chakraborty et al., 2025). For instance, in the autonomous driving context, hallucinated object detection, such as mistakenly identifying a non-existent pedestrian, could lead to severe safety incidents like abrupt stops and potential collisions. Although substantial research has addressed hallucination in LLMs (Song et al., 2024b; Yao et al., 2024; Huang et al., 2025, 2024; Zhou et al., 2024a; Arteaga et al., 2024) and in FM for autonomous driving (Dona et al., 2024; Fan et al., 2024), the underlying triggers and effective detection methods remain unclear (Xu et al., 2023). Potential research directions include leveraging multi-modal FMs to ground language with visual information (Chakraborty et al., 2025) and developing mitigation strategies that preserve the models’ generation and reasoning capabilities.
3.1.2. Navigation and Planning
FMs integrate natural language into navigation and planning by converting textual instructions into spatial representations. Systems like Talk to the Vehicle (Sriram et al., 2019) and Ground then Navigate (Jain et al., 2023) generate waypoints and trajectories from multi-modal inputs. ALT-Pilot (Omama et al., 2023) enhances planning with language-augmented maps using CLIP (Radford et al., 2021), while GPT-Driver (Mao et al., 2023) and DriveVLM (Tian et al., 2024a) support predictive planning and reasoning.
3.1.3. Decision-Making and Control
FMs improve decision-making and control by translating scene understanding into safe actions. LLMs in Drive as You Speak (Cui et al., 2024a) and LanguageMPC (Sha et al., 2023) process complex data for real-time decisions. Hybrid systems like BEVGPT (Wang et al., 2024d) and Driving with LLM (Chen et al., 2024) combine reasoning with traditional controls, while SurrealDriver (Jin et al., 2024b) and Drive Like a Human (Fu et al., 2024) enhance robustness through safety and memory modules.
3.1.4. End-to-End Autonomous Driving
Recent advancements in FMs have enabled the development of unified models that integrate perception, reasoning, and control into a single differentiable framework. DriveGPT4 (Xu et al., 2024b) processes sensor inputs and queries for control signals and explanations. ADAPT (Jin et al., 2023) maps video to actions and narratives, DriveMLM (Wang et al., 2023) integrates LLMs into closed-loop systems, and VLP (Pan et al., 2024a) promotes generalization with context-aware frameworks.
@@ 3.1.4. End-to-End Autonomous Driving
Challenge & Opportunity: Multi-modality Adaptation. Foundation models, particularly LLMs, and MLLMs, have demonstrated remarkable reasoning capabilities in autonomous driving tasks. However, most existing approaches heavily rely on environmental information from upstream perception modules, making them vulnerable to input errors (Gao et al., 2024). Even minor perception inaccuracies, such as inaccurate object heading estimation, can lead to catastrophic failures in decision-making (Mao et al., 2024). This highlights the critical need for research into robust adaptation methods that can better handle uncertainties and errors in perception inputs while maintaining reliable decision-making capabilities.
@@ 3.2. FM-Enabled Intelligent User Experience
Challenge & Opportunity: Intelligent User Interface and Personalization. While FMs enable more intelligent and personalized user experiences in autonomous vehicles, several challenges need to be addressed. MLLMs like GPT-4V can interpret natural language instructions to control vehicles according to user preferences. For example, Cui et al. demonstrated that LLM-based planners can respond to personalized commands such as “drive aggressively,” adjusting vehicle behavior across different speeds and risk levels (Cui et al., 2024b). However, this flexibility raises significant safety concerns. As shown in (Cui et al., 2024d), LLMs may interpret and execute potentially dangerous commands like “drive as fast as you can.” Although research has explored methods to ensure compliance with traffic rules and safety requirements (Cui et al., 2024d; Yang et al., 2024c), the vulnerability to jailbreak attacks remains a concern, particularly given the proliferation of LLM exploitation techniques (Rao et al., 2024; Zou et al., 2023; Zhu et al., 2024b; Jin et al., 2024a). Additionally, balancing real-time responsiveness with user privacy presents another significant challenge, as discussed in Section 2.3.
@@ 4. Foundation Model Application in Practice
This section explores the practical deployment of FMs in autonomous driving. We distinguish between modular integration and full adoption of FMs, showcasing their role in enhancing vehicle capabilities.
@@
Challenge: Foundation Model Alignment. As FMs become increasingly integrated into autonomous driving systems, their potential societal risks demand careful consideration. The undesired behaviors exhibited by FMs, such as hallucination, raise particular concerns in safety-critical domains like autonomous driving where they directly impact public safety. AI alignment has emerged as a potential solution, aiming to ensure AI systems behave in accordance with human intentions and values (Leike et al., 2018). Despite its critical importance for the safe deployment of FMs in autonomous driving, research in this area remains limited (Kong et al., 2024; IEEE, 2025; Abbo et al., 2024; Wolf et al., 2024). The complexity of foundation model systems, encompassing fairness, privacy, and security concerns, urgently calls for more attention and investigation into alignment strategies. Current alignment research can basically be divided into two key components: forward alignment and backward alignment (Ji et al., 2024), below are potential opportunities:
@@
  • Opportunity: Safety Benchmarks and Evaluation for Assurance (Backward Alignment). Datasets and benchmarks are crucial for safety evaluation, serving as fundamental tools for ensuring AI alignment. A key opportunity lies in developing comprehensive metrics and benchmarks for FMs to better evaluate their safety performance and ability to minimize accidents during task execution (Ji et al., 2024). Unlike traditional deep learning models, FMs can leverage general knowledge rather than actual cues to achieve unexpectedly high scores on existing metrics (Xie et al., 2025). This limitation highlights the need for more comprehensive benchmarks and metrics that can accurately assess both FM capabilities and potential deviations from intended behaviors (Amodei et al., 2016).
@@ 5. Conclusion
In this paper, we conducted a systematic exploration of integrating foundation models in autonomous driving, examining three key aspects: FM infrastructure, FM in Vehicles, and their real-world applications. For each aspect, we identified critical challenges and highlighted promising research opportunities. In conclusion, we believe that although significant challenges remain in integrating FMs into autonomous driving systems, it has shown its substantial potential for advancing the field. We hope this paper can serve as a roadmap for future research directions and accelerate the development of more capable autonomous driving systems.

### v2

@@ 4. Foundation Models in Vehicle
Challenge I: Hallucination. While FMs (i.e., LLMs and VLMs) have achieved significant advancements in autonomous driving, hallucination remains a critical challenge for their safe real-world deployment. Hallucination refers to the generation of outputs that are factually incorrect, inconsistent, or nonsensical, a phenomenon to which FMs are particularly prone (Sun et al., 2025; Chakraborty et al., 2025). Following the formal taxonomy established in recent research (Chakraborty et al., 2025), these failures can be categorized into four core characteristics: (1) Compliance, where the generation violates non-negotiable hard constraints (e.g., executing an illegal maneuver); (2) Desirability, where the output fails to meet soft constraints measured by optimization objectives like fuel efficiency or passenger comfort; (3) Relevancy, where the model introduces extraneous or off-topic details that do not belong to the driving task; and (4) Plausibility, which assesses the syntactic soundness of the output. Crucially, a hallucination may appear highly plausible and believable to a human critic while still being non-compliant or undesirable, making it a severe risk for safe deployment. Table 6 summarizes the representative hallucination metrics across various task settings, adapted from (Chakraborty et al., 2025). In the context of autonomous driving, a hallucinated object detection, such as mistakenly identifying a non-existent pedestrian, could trigger a severe safety incident like an abrupt stop or a potential collision. Although substantial research has addressed hallucination in general-purpose LLMs (Song et al., 2024b; Yao et al., 2024b; Huang et al., 2025, 2024; Zhou et al., 2024a; Arteaga et al., 2024) and in FMs for autonomous driving (Dona et al., 2024; Fan et al., 2026), the underlying triggers and effective detection methods remain largely unclear, creating vital opportunities for further research.
  • Opportunity: Hallucination Detection and Mitigation. While the underlying mechanism of hallucination is currently unclear, the first opportunity lies in the detection and mitigation of hallucination. To combat hallucination, research efforts have been made by both the AI and SE communities. For instance, Yang et al. (Yang et al., 2025a) utilise metamorphic relations to detect hallucinations in LLMs. Future research could focus on developing detection and mitigation strategies that reduce hallucinations without hurting the model’s performance.
  • Opportunity: Multi-modal Grounding and Verification. One promising direction to tackle hallucination is to leverage the inherently multi-modal nature of the perception data. Research may focus on developing methods to ground an FM’s generated output (e.g., textual caption) against raw sensor information from cameras, LiDAR, or radar (Chakraborty et al., 2025). Developing robust grounding techniques would anchor the model’s outputs to physical reality, which can significantly reduce the likelihood of factually incorrect or unverified statements.
Table 6. Summary of Hallucination Metrics across Task Settings (Adapted from (Chakraborty et al., 2025))
Task Setting Compliance Metrics Desirability Metrics Relevancy Metrics Plausibility Metrics
Question-Answering Accuracy, FactScore, BERTScore, Contradictions Calibration Error, Succinctness, Prudence Cross Encoder, Perplexity ROUGE, Semantic Preciseness
Image Captioning CHAIR, POPE, CLIP Score, METEOR CIDER Human Alignment, Detailedness CHAIR, POPE ROUGE, SPICE, Perplexity
Planning Feasibility, Action Probabilities, Plan/Action Accuracy Success Rate, Clarification Rate, Unsafe Action Rate, Calibration Error, Estimated Payoff Action Probabilities Non-compliance Rate, Preciseness
@@
Challenge II: Multi-modality Adaptation. While Foundation Models (FMs), particularly Large Language and Multi-modal Large Language Models, have demonstrated remarkable reasoning capabilities in autonomous driving, their practical application is often hindered by a critical vulnerability. Most current approaches rely heavily on the processed outputs of upstream perception modules, treating them as absolute ground truth (Gao et al., 2025c). This tight coupling means the entire system is brittle; even minor perception inaccuracies, such as a slight error in an object’s heading estimation, can cascade and lead to catastrophic failures in the downstream decision-making process (Mao et al., 2024). This dependency on imperfect perception highlights a critical need for robust adaptation methods, opening up several research opportunities.
  • Opportunity: Benchmarking Robustness in Multi-Modal Fusion. As research moves towards end-to-end models, there is a pressing need for standardized benchmarks to rigorously evaluate the quality and robustness of their sensor fusion capabilities. Current evaluations often focus on overall task performance, making it difficult to isolate how well a model handles sensor noise, failure, or conflicting information. This creates a significant opportunity to develop novel evaluation protocols and challenging datasets designed specifically to probe the fusion process. The creation of these benchmarks is critical for systematically comparing different architectures and guiding the development of more reliable and truly robust multimodal systems.
  • Opportunity: End-to-End Multi-Modal Fusion. A promising research direction is to move beyond cascaded pipelines and develop integrated, end-to-end FMs. The goal is to create architectures that can directly ingest and fuse raw, heterogeneous sensor data (e.g., camera pixels, LiDAR point clouds, radar signals). By learning directly from raw data, the model can develop its own robust internal representations, bypassing the vulnerabilities of a brittle intermediate perception layer. This approach allows the FM to learn complex cross-modal correlations, enabling it to rely on one modality to compensate for noise or errors in another.
@@ 4.1. FM-Enabled Intelligent User Experience
Challenge & Opportunity: Intelligent User Interface and Personalization. While FMs enable more intelligent and personalized user experiences in autonomous vehicles, several challenges need to be addressed. MLLMs like GPT-4V can interpret natural language instructions to control vehicles according to user preferences. For example, Cui et al. demonstrated that LLM-based planners can respond to personalized commands such as “drive aggressively,” adjusting vehicle behavior across different speeds and risk levels (Cui et al., 2024b). However, this flexibility raises significant safety concerns. As shown in (Cui et al., 2024d), LLMs may interpret and execute potentially dangerous commands like “drive as fast as you can.” Although research has explored methods to ensure compliance with traffic rules and safety requirements (Cui et al., 2024d; Yang et al., 2024c), the vulnerability to jailbreak attacks remains a concern, particularly given the proliferation of LLM exploitation techniques (Rao et al., 2024; Zou et al., 2023; Zhu et al., 2024c; Jin et al., 2024a). Additionally, balancing real-time responsiveness with user privacy presents another significant challenge, as discussed in Section 3.3.
@@ 5. Foundation Model Application in Practice
This section explores the practical deployment of FMs in autonomous driving. We distinguish between modular integration and full adoption of FMs, showcasing their role in enhancing vehicle capabilities. Besides, we also articulate their application in automating the development of autonomous driving systems, along with the potential challenges and opportunities.
@@
Challenge I: Foundation Model Alignment. As FMs become increasingly integrated into autonomous driving systems, their potential societal risks demand careful consideration. The undesired behaviors exhibited by FMs, such as hallucination, raise particular concerns in safety-critical domains like autonomous driving where they directly impact public safety. AI alignment has emerged as a potential solution, aiming to ensure AI systems behave in accordance with human intentions and values (Leike et al., 2018). Despite its critical importance for the safe deployment of FMs in autonomous driving, research in this area remains limited (Kong et al., 2024; IEEE, 2025; Abbo et al., 2024; Wolf et al., 2024). The complexity of foundation model systems, encompassing fairness, privacy, and security concerns, urgently calls for more attention and investigation into alignment strategies. Current alignment research can basically be divided into two key components: forward alignment and backward alignment (Ji et al., 2025). Below are potential opportunities:
@@
  • Opportunity: Safety Benchmarks and Evaluation for Assurance (Backward Alignment). Datasets and benchmarks are crucial for safety evaluation, serving as fundamental tools for ensuring AI alignment. A key opportunity lies in developing comprehensive metrics and benchmarks for FMs to better evaluate their safety performance and ability to minimize accidents during task execution (Ji et al., 2025). Unlike traditional deep learning models, FMs can leverage general knowledge rather than actual cues to achieve unexpectedly high scores on existing metrics (Xie et al., 2025). This limitation highlights the need for more comprehensive benchmarks and metrics that can accurately assess both FM capabilities and potential deviations from intended behaviors (Amodei et al., 2016).
  • Opportunity: Online Methods for Safety Alignment. While safety benchmarks can partially ensure alignment, online methods are still needed to guarantee safety during task execution in autonomous vehicles. However, due to the complexity of FMs and the requirement for low-latency execution, designing such methods remains challenging. As highlighted by OpenAI, alignment must embrace uncertainty while maintaining rigorous measurement (ope, [n. d.]). This limitation highlights the need for designing suitable online methods for FMs in autonomous driving systems.
Challenge II: Ethical and Human-Like Autonomous Driving. The ultimate objective of integrating FMs into autonomous driving is to enable ethical and human-like driving behavior. However, as pointed out by Tian et al. (Tian et al., 2026), existing regulatory frameworks for autonomous vehicles (e.g., ISO 26262) primarily address functional safety and cybersecurity for conventional rule-based systems, and do not fully capture the complexities introduced by foundation models; thus, these regulations may need to be revisited and updated. Moreover, transparency and interpretability in ADS have been increasingly emphasised due to their critical roles in regulatory approval and accountability. The integration of FMs may offer new opportunities for interpretability—for example, in speech-to-speech systems, researchers can inspect intermediate textual representations produced by LLMs to facilitate error analysis (Pan et al., 2026). Nevertheless, further advances are still required to systematically embed explainability into FM-driven ADS pipelines. Finally, since FMs are often adopted for reasoning and decision-making, it is crucial to test them under morally sensitive scenarios and to characterise how risks are distributed among different road users. Recent work has begun to operationalise such moral testing via simulation-based methods (e.g., metamorphic testing) to expose potentially problematic decision patterns (Tang et al., 2025). These efforts are essential for developing ethical and human-like autonomous driving.
  • Opportunity: Moral Testing and Assurance for FM-based ADS. As FMs are increasingly adopted for high-level reasoning and decision-making in ADS, a key opportunity is to establish systematic moral testing and assurance pipelines that complement functional safety validation. Unlike conventional safety requirements, moral expectations are often value-laden, context-dependent, and may not admit a single ground-truth label, making classic test oracles difficult to define. Recent work has started to operationalise moral evaluation via simulation-based metamorphic testing, where moral meta-principles are encoded as relational properties across paired scenarios to expose inconsistent or discriminatory decision patterns (Tang et al., 2025). Building on this direction, future research could develop scalable moral test suites and benchmarks, integrate scenario generation and coverage metrics tailored to morally sensitive situations.
Challenge III: Ensuring Correctness of FM-Generated Code. With the widespread application of FM-based coding assistance (i.e., GitHub Co-pilot), autonomous driving developers will inevitably use FM-generated code during the development of the autonomous driving system. As pointed out by Chen et al. (Chen et al., 2025b), LLM could misunderstand the code description and generate code without syntactic mistakes, but still defective. Furthermore, code generated by LLMs often lacks description or context and is more difficult for human developers to understand and maintain (Liu et al., 2024a), and may introduce critical issues into the codebase if not carefully verified (Molison et al., 2025).
  • Opportunity: Automated Verification for FM-generated Code. With the growing coding capabilities of LLMs, the use of LM-generated code is becoming inevitable. However, errors in such code can be more difficult for human developers to detect. Therefore, in safety-critical domains like autonomous driving, there is a pressing need to develop automated verification techniques for FM-generated code. Nouri et al. (Nouri et al., 2025) proposed an iterative loop to automatically generate, verify, and refine LLM-generated code in the context of autonomous driving. Yet, due to issues such as LLMs’ sensitivity to minor text disruptions in code descriptions or prompts, more robust automated verification techniques are still required.
6. Future Research Roadmap
The development and deployment of Foundation Models for Autonomous Driving (FM4AD) is a monumental task that requires a coordinated, multi-stage research effort. To guide the community, we propose a roadmap divided into three phases: short-term foundational work, mid-term system integration, and a long-term vision for achieving verifiable trustworthiness at scale.
6.1. Short-Term Goal
The immediate focus must be on establishing the fundamental building blocks required for robust and scalable research and development of FMs for autonomous driving. This involves addressing the most pressing challenges in data, development tools, and deployment efficiency.
  • Establishing Data Best Practices: The quality and management of data are paramount. The community’s short-term goals should be to develop unified data management frameworks to handle the massive scale of autonomous driving datasets and to create automated tools for dataset license compliance to ensure legal and ethical usage. Simultaneously, advancing techniques for bias mitigation and privacy preservation during data cleaning and curation is critical to building a fair and secure foundation. Success in this phase will be measured by the development of automated fairness testing suites that achieve a high correlation between perception-level bias metrics (e.g., AOD) and decision-level safety violations.
  • Improving the Development Toolchain: To manage the unprecedented complexity of the FM toolchain, the immediate priority is to conduct large-scale empirical analyses to identify and eliminate inefficiencies in current workflows. A parallel effort is required to develop a new generation of Quality Assurance and testing techniques specifically designed for the distributed systems (e.g., Ray) that underpin FM training, thereby improving the reliability and reproducibility of the development process.
  • Enabling Efficient and Safe Deployment: Before FMs can be widely integrated, their efficiency and safety on resource-constrained vehicle hardware must be understood and improved. A crucial short-term task is to perform a comprehensive empirical analysis of the impact of optimization techniques like quantization and pruning on model trustworthiness. This knowledge will directly inform the development of robust-by-design optimization methods, where security and reliability are co-designed with performance. Furthermore, success in this area will be measured by the ability of optimized models to maintain an inference latency within the 500ms range, a critical threshold for ensuring real-time responsiveness in complex driving reasoning tasks (Tian et al., 2026).
6.2. Mid-term Goal
With a solid foundation in place, the research focus can shift to the complex challenges of integrating FMs into the vehicle as a cohesive and reliable system. This phase emphasizes system-level reliability and architecture.
  • Enhancing System Reliability and Adaptation: The mid-term priority is to address the core reliability issues of FMs in dynamic environments, namely hallucination and multi-modality adaptation. This involves developing end-to-end multi-modal fusion models that are less vulnerable to upstream perception errors and using multi-modal grounding to verify model outputs against raw sensor data. To measure progress, creating standardized benchmarks for robustness in multi-modal fusion is essential.
  • Designing Hybrid System Architectures: An integrated FM-powered vehicle will not operate in isolation. A key mid-term goal is to architect effective edge-cloud collaboration systems. Research must deliver frameworks for intelligent task orchestration and robust, asynchronous communication schemes that can dynamically balance the low-latency needs of the vehicle (edge) with the immense computational power of the cloud.
  • Fostering a Domain-Specific Ecosystem: To accelerate innovation and democratize research, the community must address the lack of powerful, open-source models. A central mid-term objective should be the development of an open-source, domain-specific foundation model for autonomous driving that can serve as a strong baseline for academic and independent researchers.
6.3. Long-term Goal
The ultimate vision is the widespread deployment of autonomous vehicles powered by FMs that are not just capable but are verifiably safe, trustworthy, and aligned with human values. This requires tackling the most fundamental challenges from the hardware to the AI’s core objectives.
  • Establishing a Hardware Root of Trust: Trustworthiness must begin at the silicon level. The long-term grand challenge is to overcome the risks of ”black-box” hardware by developing security-by-design hardware architectures. This includes creating Trusted Execution Environments (TEEs) tailored for AI accelerators and promoting open and verifiable hardware stacks that allow for community-driven security audits, providing a transparent and secure foundation for all software.
  • Achieving Foundation Model Alignment: Ensuring that FMs behave in accordance with human intentions is perhaps the most critical long-term goal. This demands a sustained research program in AI alignment. The focus should be on developing robust methods for both forward alignment, such as enhancing human-feedback mechanisms during training, and backward alignment, by creating comprehensive safety benchmarks to evaluate model behavior. A key frontier is the design of effective online methods for safety alignment that can provide real-time assurance in a deployed vehicle.
  • Guaranteeing Code and System Correctness: As FMs increasingly contribute to writing the software that runs autonomous vehicles, verifying its correctness is a long-term imperative. The research community must work towards powerful automated verification techniques that can formally prove the safety and reliability of FM-generated code, ensuring that this powerful tool enhances, rather than compromises, the safety of autonomous systems.
@@ 7. Conclusion
In this paper, we presented a structured literature review and an initial roadmap for integrating foundation models into autonomous driving systems. We organized the discussion around three dimensions: FM infrastructure, in-vehicle integration, and practical deployment. Across these dimensions, we synthesized the state of the art, identified key challenges, and highlighted promising research opportunities, which we further distilled into short-, mid-, and long-term research goals. Although substantial technical and engineering challenges remain, FMs offer significant potential to advance the capabilities of ADSs. We hope this paper serves as a useful reference for future research and helps guide the development of safer, more reliable, and more trustworthy autonomous driving systems.
Acknowledgements.
This work was partially supported by the National Natural Science Foundation of China: No. 52408039, JST CRONOS Grant (No. JPMJCS24K8) and JSPS KAKENHI Grant (No.JP24K02920). This research/project was also supported by the Ministry of Education, Singapore under its Academic Research Fund Tier 2 (Proposal ID: T2EP20223-0043; Project ID: MOE-000613-00). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.